7. Calculate the heat of following reaction N 2 + 3H 2 —> 2NH 3 Given the bond energies of N = N, H - H and N ¾ H bonds are 226, 104 and 93 kcal respectively. 8. When 2 moles of C 2 H 6 are completely burnt 3129 KJ of heat is liberated. Calculate the heat of formation, ΔH f for C 2 H 6 ; Δ H f for CO 2 ­ and H 2 O are -395 and -286 KJ respectively. 9. Calculate the heat of formation of ethane at 25°C. The bond enthalpies for H - H, C - C and C -H bonds are 104.2 kcal, 80 kcal and 99.5 kcal respectively. Heat of vaporization of carbon is 171.7 kcal. 10. Define the following terms: (a) Internal energy (b) Endothermic reaction (c) Hess law (d) Calorific value 11. 5 mole of an ideal gas expand isothermally & reversibly from a pressure of 10 atm to 2 atm at 300 K. What is the largest mass which can be lifted through a height of 1 mitre in this expansion? 12. The equilibrium constant for the reaction: CO 2 (g) + H 2 (g) ——> CO(g) + H 2 O at 298 K is 73. Calculate the value of the standard free energy change (R = 8.314 JK - 1 mol - 1 ). 13. An insulated container contains 1 mole of a liquid molar volumes 100 ml at 1 bar. When liquid is steeply passed to 100 bar, volume decrease t0 99 ml, find Δ H & ΔU for the process.


7.   Calculate the heat of following reaction


 

      N2 + 3H2 —> 2NH3


 

      Given the bond energies of N = N, H - H and N ¾ H bonds are 226, 104 and 93 kcal respectively.


 

8.   When 2 moles of C2H6 are completely burnt 3129 KJ of heat is liberated. Calculate the heat of formation, ΔHf for C2H6ΔHf for CO2­ and H2O are -395 and -286 KJ respectively.


 

9.   Calculate the heat of formation of ethane at 25°C. The bond enthalpies for H - H, C - C and C -H bonds are 104.2 kcal, 80 kcal and 99.5 kcal respectively. Heat of vaporization of carbon is 171.7 kcal. 


 

10. Define the following terms:


 

      (a) Internal energy            (b) Endothermic reaction


 

      (c) Hess law                    (d) Calorific value


 

11.5 mole of an ideal gas expand isothermally & reversibly from a pressure of 10 atm to 2 atm at 300 K. What is the largest mass which can be lifted through a height of 1 mitre in this expansion?


 

12.The equilibrium constant for the reaction:


 

      CO2(g) + H2(g) ——> CO(g) + H2O at 298 K is 73. Calculate the value of the standard free energy change (R = 8.314 JK-1mol-1).


 

13.An insulated container contains 1 mole of a liquid molar volumes 100 ml at 1 bar. When liquid is steeply passed to 100 bar, volume decrease t0 99 ml, find ΔH & ΔU for the process.


 

Grade:

1 Answers

maurya ksk
11 years ago

7.  1*226+3*104=638kcal gives total energy of the reactants

     in NH3 3 N-H bonds are present 

     So, 3*93=279kcal

     but there are 2 NH3 molecules

     therefore

     2*279=558kcal

     net heat=558-638=-80kacl

     which shows reaction is exothermic.